* EQ Southeast of LA around 45 minutes ago. Callum and I felt it.
* Koji and I came in to recover. MC suspensions had been mis-aligned. ETMs both tripped their watchdogs.
* As before, the ETMX was stuck in its cage and the UR & LR OSEMs were reading zero V.
* We moved the MC sus back to their OSEM values from 2 hours ago. Koji aligned everything else by just using his chee.
* To shake the ETMX loose, I tried a different tactic than the "Great Balls of Fire". I started giving it 20k steps through the ASCYAW filterbank (with ramping OFF). I used the green light in the X arm video to look at the swinging. Using this as a readback I pumped the OFFSET button on ASCYAW to resonantly swing up the yaw motion. I had to turn the watchdog thresh up to 2000 temporarily. After a couple minutes the ETMX was free.
* We then used the bias sliders to steer it back onto the OL center (which Q nicely lined up for us recently) and then X arm locked in green right away.
Fri Mar 28 22:38:04 2014: We've just ridden through the 5th aftershock. None of the aftershocks have tripped the watchdogs but they break the IMC lock.
Local earthquake activity is up. Our suspensions are holding well. ETMX and ETMY sus damping restored.